how many liters of the 30% acid solution should be added to 40 liters of 10% solution to get a 25% solution? (please state the meaning of any variables you use, set up an algebraic equation, and use it to solve the problem)
Answer by lwsshak3(11628)

how many liters of the 30% acid solution should be added to 40 liters of 10% solution to get a 25% solution? (please state the meaning of any variables you use, set up an algebraic equation, and use it to solve the problem)
***
let x=amount of the 30% acid solution should be added to 40 liters of 10% solution
amount of final 25% solution=(40+x)
sum of acid in the two solutions=amt of acid in final solution
30%x+10%40=25%(40+x)
.30x+4=10+.25x
.05x=6
x=120
amount of the 30% acid solution should be added to 40 liters of 10% solution=120 liters